the 5 g of zinc powder used is an excess. Calculate the mass of zinc that reacts exactly with 50.0cm3 of 1.00 mol dm


1
Expert's answer
2022-01-25T02:55:05-0500

Zn + CuSO4 = ZnSO4 + Cu

According to the equation, n (Zn) = n (CuSO4)

CM = n / V

n (CuSO4) = CM x V = 1.00 x 0.0500 = 0.05 mol

n = m / M

M (Zn) = 65.38 g/mol

m (Zn) = n x M = 0.05 x 65.38 = 3.269 g